# FactReasoner: A Probabilistic Approach to Long-Form Factuality Assessment for Large Language Models

### v2

@@ Abstract
Large language models (LLMs) have achieved remarkable success in generative tasks, yet they often fall short in ensuring the factual accuracy of their outputs this limiting their reliability in real-world applications where correctness is critical. In this paper, we present FactReasoner, a novel factuality assessment framework that employs probabilistic reasoning to evaluate the truthfulness of long-form generated responses. FactReasoner decomposes a response into atomic units, retrieves relevant contextual information from external knowledge sources, and models the logical relationships (e.g., entailment, contradiction) between these units and their contexts using probabilistic encodings. It then estimates the posterior probability that each atomic unit is supported by the retrieved evidence. Our experiments on both labeled and unlabeled benchmark datasets demonstrate that FactReasoner often outperforms state-of-the-art prompt-based methods in terms of factual precision and recall.
@@
Radu Marinescu1, Debarun Bhattacharjya1, Junkyu Lee1, Tigran Tchrakian1, Javier Carnerero Cano1, Yufang Hou1,2, Elizabeth Daly1, Alessandra Pascale1, 1IBM Research, 2IT:U - Interdisciplinary Transformation University Austria, Correspondence: radu.marinescu@ie.ibm.com
@@ 1 Introduction
Context Relevance Across Atoms: In multi-atom responses, contexts retrieved for one atom can be relevant – either supportive or contradictory – to another atom. Prompt-based methods struggle with this, as they require saturating the model’s context window with all of the retrieved information. FactReasoner overcomes this limitation using a compact probabilistic representation (i.e., a graphical model) of the relationships between all atoms in the response and all of the retrieved contexts.
Handling Conflicting Contexts: Sometimes, contexts retrieved for different atoms may contradict each other. FactReasoner can leverage these contradictions effectively and in a principled manner by reasoning over their probabilistic encodings which often leads to improved performance.
Leveraging LLM Strengths in NLI Tasks: LLMs excel at natural language inference tasks such as entailment and contradiction. FactReasoner builds on this strength by framing factuality assessment as a composition of these simpler tasks.
@@ 2 Background
Graphical Models. A graphical model is a tuple $\mathcal{M}=\langle\mathbf{X},\mathbf{D},\mathbf{F}\rangle$, where $\mathbf{X}=\{X_{1},\ldots,X_{n}\}$ is a set of variables, $\mathbf{D}=\{D_{1},\ldots,D_{n}\}$ is the set of their finite domains of values and $\mathbf{F}=\{f_{1},\ldots,f_{m}\}$ is a set of discrete positive real-valued functions. Each function $f_{i}$ (also called factor) is defined on a subset of variables $\mathbf{S}_{i}\subseteq\mathbf{X}$ called its scope and denoted by $vars(f_{i})$. The model $\mathcal{M}$ defines a factorized probability distribution on $\mathbf{X}$: $P(\mathbf{x})=\frac{1}{Z}\prod_{j=1}^{m}f_{j}(\mathbf{x})$ where $Z=\sum_{\mathbf{x}\in\Omega(\mathbf{X})}\prod_{j=1}^{m}f_{j}(\mathbf{x})$ is the normalization constant $Z$ also known as the partition function and $\Omega(\mathbf{X})$ denotes the Cartesian product of the variables domains Koller and Friedman (2009).
A common inference task over graphical models is to compute the posterior marginal distributions over all variables. Namely, for each variable $X_{i}\in\mathbf{X}$ and domain value $x_{i}\in D_{i}$, compute: $P(x_{i})=\sum_{\mathbf{x}\in\Omega(\mathbf{X})}\delta_{x_{i}}(\mathbf{x})\cdot P(\mathbf{x})$, where $\delta_{x_{i}}(\mathbf{x})$ is 111 if $X_{i}$ is assigned $x_{i}$ in $\mathbf{x}$ and 0 otherwise.
@@
Long-Form Factuality. Let $y$ be the long-form response generated by an LLM to a query $x$. Following prior work Min et al. (2023); Song et al. (2024); Wei et al. (2024), we assume that $y$ can be decomposed into a set of $n$ atomic units (or atoms) that can be either true or false, denoted by $\mathcal{A}_{y}=\{a_{1},a_{2},\ldots a_{n}\}$. An atomic unit $a_{i}\in\mathcal{A}_{y}$ is defined as a short sentence conveying one piece of information. Furthermore, given an external knowledge source $\mathcal{C}$ [^1] For example, $\mathcal{C}$ could be Wikipedia, Google Search, or a collection of documents embedded into a vector database., we say that an atomic unit $a_{i}\in\mathcal{A}_{y}$ is supported by $\mathcal{C}$ if there exists at least one piece of information in $\mathcal{C}$ (e.g., a passage) called a context that undebatably supports $a_{i}$. Otherwise, we say that the atomic unit is not supported. The factual precision $Pr(y)$ of the response $y$ with respect to a knowledge source $\mathcal{C}$ is defined as: $Pr(y)=\frac{S(y)}{|\mathcal{A}_{y}|}$, where $S(y)=\sum_{i=1}^{n}\mathbb{I}[a_{i}\textrm{ is supported by }\mathcal{C}]$ is the number of supported atomic units. Furthermore, the notion of factual recall up to the $K$-th supported atomic unit denoted by $R_{K}(y)$ is given by: $R_{K}(y)=\min(\frac{S(y)}{K},1)$. Finally, an $F_{1}$ measure for long-form factuality denoted by $F1@K$ can be defined as: $F_{1}@K(y)=\frac{2\cdot Pr(y)\cdot R_{K}(y)}{Pr(y)+R_{K}(y)}$ if $S(y)>0$, and 0 otherwise Wei et al. (2024).
@@ 3.1 A Graphical Models Based Approach
Variables. We associate each atom $a_{i}\in\mathcal{A}_{y}$ and context $c_{j}\in\mathcal{C}_{y}$ with a bi-valued variable denoted by either $A_{i}$ (for atoms) or $C_{j}$ (for contexts). Therefore, we have that $\mathbf{X}=\mathbf{X}_{a}\cup\mathbf{X}_{c}$ where $\mathbf{X}_{a}=\{A_{1},\ldots,A_{n}\}$ and $\mathbf{X}_{c}=\{C_{1},\ldots,C_{m}\}$, respectively. The domains of the variables contain the values true and false indicating whether the corresponding atom or context is true or false. For simplicity, we use $a_{i}$ and $\neg a_{i}$ (resp. $c_{j}$ and $\neg c_{j}$) to denote the value assignments $A_{i}=true$ and $A_{i}=false$ (resp. $C_{j}=true$ and $C_{j}=false$).
Priors. For each variable $A_{i}\in\mathbf{X}_{a}$ (resp. $C_{j}\in\mathbf{X}_{c}$) we consider a unary factor denoted by $f(A_{i})$ (resp. $f(C_{j})$) representing the prior belief about the truthfulness of the corresponding atom (resp. context). Since we make no assumptions about the response, we set $f(a_{i})=0.5$ and $f(\neg a_{i})=0.5$, respectively. In contrast, the external knowledge source $\mathcal{C}$ is assumed to be reliable and therefore the retrieved contexts have high probability of being true (e.g., $f(c_{j})=0.99$). Note that if a context is retrieved from a less reliable source then its prior probability can be set to a smaller value.
Relationships. In addition, we also consider binary factors denoted by $f(A_{i},C_{j})$ and $f(C_{j},C_{k})$, defined on atom-context variable pairs as well as pairs of context variables. These factors are probabilistic representations of the logical relationships between the natural language utterances corresponding to the context and atom variables. For our purpose, we use a relation model $p_{\theta}(\cdot|t,t^{\prime})$ to predict the most likely logical relationship between an ordered pair of natural language utterances from the choices {none, entail, contradict, equivalence} [^2] The “equivalence” relationship is formed if entailment is predicted for both orderings of the utterances. The “none” relationship corresponds to neutrality meaning that the two utterances are not related to each other.. The relation model can be any pre-trained BERT or LLM Liu et al. (2019); Touvron et al. (2023).
@@
[…]
@@
$\mathbf{F}$ =$=\{f(C_{j},A_{i})~|~A_{i}\in\mathbf{X}_{a},C_{j}\in\mathbf{X}_{c}\}$
@@
[…]
@@ 3.3 The FactReasoner Pipeline
FactReasoner 1 (FR1). In this case, for each atom variable $A_{i}$ up to $k$ most relevant contexts $\{C_{1}^{i},...,C_{k}^{i}\}$ are retrieved and only the relationships between each atom $A_{i}$ and its corresponding contexts are considered, namely only the factors $f(A_{i},C_{j}^{i})$ are created (where $j=1..k$).
FactReasoner 2 (FR2). This version also retrieves up to $k$ contexts for each atom $A_{i}$, but it subsequently removes any duplicated contexts, thus resulting in $m$ unique contexts denoted by $\{C_{1},...,C_{m}\}$. It then considers the relationships between each atom $A_{i}$ and all $m$ contexts, creating the factors $f(A_{i},C_{j})$, where $j=1..m$.
FactReasoner 3 (FR3). We consider the same contexts $\{C_{1},...,C_{m}\}$ as in FR2, but in addition to the atom-context relationships we also consider the context-context relationships. Thus, we create the factors $f(A_{i},C_{j})$ and $f(C_{j},C_{k})$, where $j=1..m$, $k=1..m$ and $j\neq k$, respectively.
@@ 4 Experiments
In this section, we empirically evaluate the FactReasoner assessor for long-form factuality and compare it against state-of-the-art approaches on labeled and unlabeled datasets. Although the FactReasoner pipeline stages can be instantiated with different LLMs, in our implementation we use the same LLM throughout the entire pipeline and focus our empirical evaluation on the Evaluator stage (i.e., factuality assessment).
Baseline Assessors. For our purpose, we consider the following state-of-the-art prompt-based long-form factuality assessors: FactScore (FS) Min et al. (2023), FactVerify (FV) Bayat et al. (2025) and VeriScore (VS) Song et al. (2024). FactScore is one of the first assessor that prompts an LLM to assess whether an atomic unit of the response is supported or not by a set of contexts relevant to the atom which are retrieved from an external knowledge source such as Wikipedia. FactVerify and VeriScore are more recent refinements of FactScore’s original prompt that can accommodate other external knowledge sources such as Google Search results and enable the LLM’s reasoning capabilities to evaluate the relationships between an atom and its relevant contexts. Unlike FactScore, the latter can label the atoms as supported, contradicted and undecided, respectively. In our experiments, we instantiated the competing assessors including the FactReasoner variants with open-source LLMs belonging to the IBM Granite [^3] https://huggingface.co/ibm-granite, Meta LLaMA [^4] https://huggingface.co/meta-llama and MistralAI Mixtral [^5] https://huggingface.co/mistralai families, namely: granite-3.0-8b-instruct, llama-3.1-70b-instruct, and mixtral-8x22b-instruct, respectively. All our LLMs are hosted remotely on compute nodes with A100 80GB GPUs and accessed via litellm APIs capable of serving 1500 prompts per second.
Datasets. We experimented with the following datasets: Biographies (Bio) Min et al. (2023), AskHistorians (AskH) Xu et al. (2023), ELI5 Xu et al. (2023), FreshBooks (Books) Song et al. (2024), and LongFact-Objects (LFObj) Wei et al. (2024). These datasets have been widely adopted in prior work and are considered representative benchmarks for long-form factuality assessment, as they encompass a diverse range of topics and tasks, including creative writing, history, astronomy, chemistry, and more.
@@
Measures of Performance. For each dataset $\mathcal{D}$ and each competing assessor, we report the factual precision ($Pr$) and the $F_{1}@K$ score, averaged over all prompts in $\mathcal{D}$. If $\mathcal{D}$ includes annotated atomic units (i.e., ground truth labels), we additionally report the standard $F_{1}$ score and the mean absolute error (MAE), defined as:
@@
External Knowledge Sources. We consider two external knowledge sources: Wikipedia and Google Search results. For a given atom, the top $k$ results are retrieved as contexts either from wikipedia.org using the Wikipedia retriever available from LangChain [^6] https://python.langchain.com, or from google.com using the Serper API [^7] https://serper.dev. In both cases, a context is a tuple $(t,l,s,d)$, where $t$ is the title of the wiki/web-page, $l$ is the link, $s$ is a short text snippet or summary and $d$ is the content retrieved from $l$ (but capped at max 4000 characters). We used $k=3$ for the Wikipedia retriever and $k=5$ for the Google Search results Min et al. (2023); Wei et al. (2024).

### v3

@@ Abstract
Large language models (LLMs) have achieved remarkable success in generative tasks, yet they often fall short in ensuring the factual accuracy of their outputs, thus limiting their reliability in real-world applications where correctness is critical. In this paper, we present FactReasoner, a novel neuro-symbolic based factuality assessment framework that employs probabilistic reasoning to evaluate the truthfulness of long-form generated responses. FactReasoner decomposes a response into atomic units, retrieves relevant contextual information from external knowledge sources, and models the logical relationships (e.g., entailment, contradiction) between these units and their contexts using probabilistic encodings. It then estimates the posterior probability that each atomic unit is supported by the retrieved evidence. Our experiments on both labeled and unlabeled benchmark datasets demonstrate that FactReasoner often outperforms state-of-the-art prompt-based methods in terms of factual precision and recall. Our open-source implementation is publicly available at: https://github.com/IBM/FactReasoner.
@@
Radu Marinescu1, Debarun Bhattacharjya1, Junkyu Lee1, Tigran Tchrakian1, Javier Carnerero Cano1, Yufang Hou1,2, Elizabeth Daly1, Alessandra Pascale1 1IBM Research, 2IT:U - Interdisciplinary Transformation University Austria Correspondence: radu.marinescu@ie.ibm.com
@@ 1 Introduction
Context Relevance Across Atoms
In multi-atom responses, contexts retrieved for one atom can be relevant – either supportive or contradictory – to another atom. Prompt-based methods struggle with this, as they require saturating the model’s context window with all of the retrieved information. FactReasoner overcomes this limitation using a compact probabilistic representation (i.e., a graphical model) of the relationships between all atoms in the response and all of the retrieved contexts.
Handling Conflicting Contexts
Sometimes, contexts retrieved for different atoms may contradict each other. FactReasoner can leverage these contradictions effectively and in a principled manner by reasoning over their probabilistic encodings which often leads to improved performance.
Leveraging LLM Strengths in NLI Tasks
LLMs excel at natural language inference tasks such as entailment and contradiction. FactReasoner builds on this strength by framing factuality assessment as a composition of these simpler tasks.
@@ 2.1 Graphical Models
A graphical model is a tuple $\mathcal{M}=\langle\mathbf{X},\mathbf{D},\mathbf{F}\rangle$, where $\mathbf{X}=\{X_{1},\ldots,X_{n}\}$ is a set of variables, $\mathbf{D}=\{D_{1},\ldots,D_{n}\}$ is the set of their finite domains of values and $\mathbf{F}=\{f_{1},\ldots,f_{m}\}$ is a set of discrete positive real-valued functions. Each function $f_{i}$ (also called factor) is defined on a subset of variables $\mathbf{S}_{i}\subseteq\mathbf{X}$ called its scope and denoted by $vars(f_{i})$. The function scopes of a model $\mathcal{M}$ define a primal graph whose vertices are the variables and its edges connect any two variables that appear in the scope of the same function.
The model $\mathcal{M}$ defines a factorized probability distribution on $\mathbf{X}$:
$P(\mathbf{x})$ $=\frac{1}{Z}\prod_{j=1}^{m}f_{j}(\mathbf{x})$ (1)
where $Z=\sum_{\mathbf{x}\in\Omega(\mathbf{X})}\prod_{j=1}^{m}f_{j}(\mathbf{x})$ is the normalization constant $Z$ also known as the partition function and $\Omega(\mathbf{X})$ denotes the Cartesian product of the variables domains Koller and Friedman (2009).
where $\delta_{x_{i}}(\mathbf{x})$ is $1$ if $X_{i}$ is assigned $x_{i}$ in $\mathbf{x}$ and 0 otherwise Koller and Friedman (2009).
Equation 2 can be solved using any probabilistic inference algorithm for graphical models, such as variable elimination Dechter (2003), belief propagation Pearl (1988); Dechter et al. (2002); Yedidia et al. (2005), sampling Bidyuk and Dechter (2003); Gogate and Dechter (2007), search Mateescu and Dechter (2005); Dechter and Mateescu (2007); Marinescu and Dechter (2009), partitioning based approximations Dechter and Rish (2003), or variational inference Wainwright and Jordan (2008); Liu and Ihler (2011).
2.2 Long-Form Factuality
Let $y$ be the long-form response generated by an LLM to a query $x$. Following prior work Min et al. (2023); Song et al. (2024); Wei et al. (2024), we assume that $y$ can be decomposed into a set of $n$ atomic units (or atoms) that can be either true or false, denoted by $\mathcal{A}_{y}=\{a_{1},a_{2},\ldots a_{n}\}$.
@@ 2.2 Long-Form Factuality
An atomic unit $a_{i}\in\mathcal{A}_{y}$ is defined as a short sentence conveying one piece of information. Furthermore, given an external knowledge source $\mathcal{C}$ [^1] For example, $\mathcal{C}$ could be Wikipedia, Google Search, or a collection of documents embedded into a vector database., we say that an atomic unit $a_{i}\in\mathcal{A}_{y}$ is supported by $\mathcal{C}$ if there exists at least one piece of information in $\mathcal{C}$ (e.g., a passage) called a context that undebatably supports $a_{i}$. Otherwise, we say that the atomic unit is not supported.
The factual precision $Pr(y)$ of the response $y$ with respect to a knowledge source $\mathcal{C}$ is defined as:
$Pr(y)$ $=\frac{S(y)}{|\mathcal{A}_{y}|}$ (3)
where $S(y)=\sum_{i=1}^{n}\mathbb{I}[a_{i}\textrm{ is supported by }\mathcal{C}]$ is the number of supported atomic units.
Furthermore, the notion of factual recall up to the $K$-th supported atomic unit denoted by $R_{K}(y)$ can be defined as follows:
$R_{K}(y)=\min(\frac{S(y)}{K},1)$ (4)
Combining Equations 3 and 4 yields an $F_{1}$ measure for factuality denoted $F1@K$ as follows:
$F_{1}@K(y)$ $=\left\{\begin{aligned} &\frac{2\cdot Pr(y)\cdot R_{K}(y)}{Pr(y)+R_{K}(y)},&S(y)>0\\ &0,&S(y)=0\end{aligned}\right$.
@@ 3.1 A Graphical Models Based Approach
Variables.
We associate each atom $a_{i}\in\mathcal{A}_{y}$ and context $c_{j}\in\mathcal{C}_{y}$ with a bi-valued variable denoted by either $A_{i}$ (for atoms) or $C_{j}$ (for contexts). Therefore, we have that $\mathbf{X}=\mathbf{X}_{a}\cup\mathbf{X}_{c}$ where $\mathbf{X}_{a}=\{A_{1},\ldots,A_{n}\}$ and $\mathbf{X}_{c}=\{C_{1},\ldots,C_{m}\}$, respectively. The domains of the variables contain the values true and false indicating whether the corresponding atom or context is true or false. For simplicity, we use $a_{i}$ and $\neg a_{i}$ (resp. $c_{j}$ and $\neg c_{j}$) to denote the value assignments $A_{i}=true$ and $A_{i}=false$ (resp. $C_{j}=true$ and $C_{j}=false$).
Priors.
For each variable $A_{i}\in\mathbf{X}_{a}$ (resp. $C_{j}\in\mathbf{X}_{c}$) we consider a unary factor denoted by $f(A_{i})$ (resp. $f(C_{j})$) representing the prior belief about the truthfulness of the corresponding atom (resp. context). Since we make no assumptions about the response, we set $f(a_{i})=0.5$ and $f(\neg a_{i})=0.5$, respectively. In contrast, the external knowledge source $\mathcal{C}$ is assumed to be reliable and therefore the retrieved contexts have high probability of being true (e.g., $f(c_{j})=0.99$). Note that if a context is retrieved from a less reliable source then its prior probability can be set to a smaller value.
Relationships.
In addition, we also consider binary factors denoted by $f(A_{i},C_{j})$ and $f(C_{j},C_{k})$, defined on atom-context variable pairs as well as pairs of context variables. These factors are probabilistic representations of the logical relationships between the natural language utterances corresponding to the context and atom variables. For our purpose, we use a relation model $p_{\theta}(\cdot|t,t^{\prime})$ to predict the most likely logical relationship between an ordered pair of natural language utterances from the choices {none, entail, contradict, equivalence} [^2] The “equivalence” relationship is formed if entailment is predicted for both orderings of the utterances. The “none” relationship corresponds to neutrality meaning that the two utterances are not related to each other.. The relation model can be any pre-trained BERT or LLM Liu et al. (2019); Touvron et al. (2023).
@@
For instance, if $r^{*}$ corresponds to entailment and $(X,Y)$ is a context-atom pair then the context supports the atom. Alternatively, if $r^{*}$ is a contradiction for the same $(X,Y)$ pair then the context contradicts the atom. Finally, for BERT-based relation models, the probability $p^{*}$ is given together with the predicted relationship $r^{*}$, whereas for instructed LLM-based relation models we can obtain $p^{*}$ by applying any uncertainty quantification (UQ) method Lin et al. (2024); Gao et al. (2024). We use a simple white-box UQ method that calculates $p^{*}$ using the logits of the “entailment" or “contradiction" tokens produced by the model. In our experiments, we use LLM-based relation models.
@@
$\mathbf{F}$ $=\{f(C_{j},A_{i})~|~A_{i}\in\mathbf{X}_{a},C_{j}\in\mathbf{X}_{c}\}$
@@
where we consider $r^{*}\in$ $\{\textrm{entail}$, $\textrm{contradict}\}$ for the context-atom pairs, and $r^{*}\in$ $\{\textrm{entail}$, contradict, $\textrm{equivalence}\}$ for the context pairs, respectively.
@@ 3.3 The FactReasoner Pipeline
FactReasoner 1 (FR1).
In this case, for each atom variable $A_{i}$ up to $k$ most relevant contexts $\{C_{1}^{i},...,C_{k}^{i}\}$ are retrieved and only the relationships between each atom $A_{i}$ and its corresponding contexts are considered, namely only the factors $f(A_{i},C_{j}^{i})$ are created (where $j=1..k$).
FactReasoner 2 (FR2).
This version also retrieves up to $k$ contexts for each atom $A_{i}$, but it subsequently removes any duplicated contexts, thus resulting in $m$ unique contexts denoted by $\{C_{1},...,C_{m}\}$. It then considers the relationships between each atom $A_{i}$ and all $m$ contexts, creating the factors $f(A_{i},C_{j})$, where $j=1..m$.
FactReasoner 3 (FR3).
We consider the same contexts $\{C_{1},...,C_{m}\}$ as in FR2, but in addition to the atom-context relationships we also consider the context-context relationships. Thus, we create the factors $f(A_{i},C_{j})$ and $f(C_{j},C_{k})$, where $j=1..m$, $k=1..m$ and $j\neq k$, respectively.
@@ 4 Experiments
We empirically evaluate the FactReasoner assessor for long-form factuality and compare it against state-of-the-art approaches on labeled and unlabeled datasets. Although the FactReasoner pipeline stages can be instantiated with different LLMs, in our implementation we use the same LLM throughout the entire pipeline and focus our empirical evaluation on the Evaluator stage (i.e., factuality assessment). Furthermore, our open-source Python implementation is publicly available at: https://github.com/IBM/FactReasoner.
Baseline Assessors.
For our purpose, we consider the following state-of-the-art prompt-based long-form factuality assessors: FactScore (FS) Min et al. (2023), FactVerify (FV) Bayat et al. (2025) and VeriScore (VS) Song et al. (2024). FactScore is one of the first assessor that prompts an LLM to assess whether an atomic unit of the response is supported or not by a set of contexts relevant to the atom which are retrieved from an external knowledge source such as Wikipedia. FactVerify and VeriScore are more recent refinements of FactScore’s original prompt that can accommodate other external knowledge sources such as Google Search results and enable the LLM’s reasoning capabilities to evaluate the relationships between an atom and its relevant contexts. Unlike FactScore, the latter can label the atoms as supported, contradicted and undecided, respectively. In our experiments, we instantiated the competing assessors including the FactReasoner variants with open-source LLMs belonging to the IBM Granite [^3] https://huggingface.co/ibm-granite, Meta LLaMA [^4] https://huggingface.co/meta-llama and MistralAI Mixtral [^5] https://huggingface.co/mistralai families, namely: granite-3.0-8b-instruct, llama-3.1-70b-instruct, and mixtral-8x22b-instruct, respectively. All our LLMs are hosted remotely on compute nodes with A100 80GB GPUs and accessed via litellm APIs capable of serving 1500 prompts per second.
Datasets.
@@
Measures of Performance.
@@
External Knowledge Sources.
We consider two external knowledge sources: Wikipedia and Google Search results. For a given atom, the top $k$ results are retrieved as contexts either from wikipedia.org using the Wikipedia retriever available from LangChain [^6] https://python.langchain.com, or from google.com using the Serper API [^7] https://serper.dev. In both cases, a context is a tuple $(t,l,s,d)$, where $t$ is the title of the wiki/web-page, $l$ is the link, $s$ is a short text snippet or summary and $d$ is the content retrieved from $l$ (but capped at max 4000 characters). We used $k=3$ for the Wikipedia retriever and $k=5$ for the Google Search results Min et al. (2023); Wei et al. (2024).